Monitoring - We believe this issue is now resolved however are continuing to monitor the traffic.
Mar 25, 15:21 UTC
Investigating - We are currently investigating a DDoS attack against our Coventry site which may be causing increased latency and packet loss on some routes. Further updates to follow within the next ten minutes.
Mar 25, 15:15 UTC
Coventry DH1 Operational
90 days ago
100.0 % uptime
Today
Coventry DH2 Operational
90 days ago
100.0 % uptime
Today
Telehouse East (THE) Operational
90 days ago
100.0 % uptime
Today
Telehouse North (THN) Operational
90 days ago
99.99 % uptime
Today
Telehouse West (THW) Operational
90 days ago
100.0 % uptime
Today
Sovereign House (SOV) Operational
90 days ago
100.0 % uptime
Today
Pulsant (Maidenhead) 1 Operational
90 days ago
99.98 % uptime
Today
Pulsant (Maidenhead) 2 Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
had a major outage
had a partial outage
Scheduled Maintenance
COVID-19 Coronavirus Outbreak Mar 20, 16:00 - Apr 10, 12:00 UTC
In light of the ongoing Coronavirus situation, UK Servers has implemented several precautionary procedures that reflect the company’s focus on the health and safety of our employees, customers and partners; the continuity of our services

UK Servers are fully committed to keeping all services online with no planned interruption to services.
Our staff have a number of procedures in place to ensure they are able to work from different locations if so required.

We have taken steps to ensure that support staff will be available 24/7 as they are now, however if staff numbers do reduce due to people being off ill, there may be a delayed response to support requests.

In addition, non-urgent requests such as upgrades, new orders and accounts queries may also see a delayed response as we prioritise workload.

If there are any changes to the current situation, UKServers will issue a further update communicating this to customers.

Currently our data centre sites remain accessible to all customers, however, in line with government recommendations, we encourage visits for essential works only and access to the datafloor is limited for non-essential activities.
Posted on Mar 20, 15:24 UTC
Past Incidents
Mar 31, 2020

No incidents reported today.

Mar 30, 2020

No incidents reported.

Mar 29, 2020

No incidents reported.

Mar 28, 2020

No incidents reported.

Mar 27, 2020

No incidents reported.

Mar 26, 2020

No incidents reported.

Mar 24, 2020

No incidents reported.

Mar 23, 2020

No incidents reported.

Mar 22, 2020

No incidents reported.

Mar 21, 2020

No incidents reported.

Mar 20, 2020
Resolved - This fibre link has now been fully restored, traffic is being sent over this link as normal and we do not expect any further issues.
Mar 20, 16:23 UTC
Update - This issue remains ongoing, we are currently routing traffic via alternative paths, this provider is currently waiting for another engineer to arrive onsite to re-splice this fibre at which point we expect the issue will be resolved. A further update has been promised by the provider by 13:00.
Mar 20, 12:27 UTC
Investigating - Our monitoring has shown these links have gone offline again, we have contacted the fibre provider and are awaiting more information.
Mar 20, 09:17 UTC
Update - Our service provider has given us the all clear and confirmed that these circuits have been re-spliced and we are now sending traffic over these links again. We will continue to monitor these circuits closely however the network is no longer deemed "at risk".
Mar 20, 08:53 UTC
Monitoring - We are now seeing service restored on these links, we are awaiting confirmation that the issue has been resolved before resuming traffic over these fibres.
Mar 20, 08:45 UTC
Update - Our service provider engineers have identified the fibre fault and are currently awaiting a splicer to arrive to repair the broken fibre. The splicer is expected to be onsite at 08:00, we expect to have another update from the service provider at 09:00.
Mar 20, 07:38 UTC
Identified - Our service provider has informed us that engineering teams on the ground are working to isolate this fault and implement a fix plan, we expect a further service update from them at 06:45 AM.
Mar 20, 05:58 UTC
Update - A number of wavelengths between our Coventry Datacentre and London POPs have failed.
We are working with our connectivity providers to ensure this is resolved as soon as possible, meanwhile traffic has been balanced in an attempt to minimise disruption to service during this time.
Mar 20, 03:29 UTC
Investigating - We are investigating a connection issue between our Coventry datacentre and London core network
Some customers may see packet loss or higher latency as a result of this.

Further updates to follow.
Mar 20, 03:08 UTC
Mar 19, 2020

No incidents reported.

Mar 18, 2020

No incidents reported.

Mar 17, 2020
Resolved - This incident has been resolved.
Mar 17, 10:34 UTC
Investigating - As of 23:41:16 GMT our monitoring has reported a loss in one of the wavelengths connecting the Maidenhead datacentre site to our Telehouse East London POP.

Due to the redundant nature of our network customers should not be experiencing any issues related to this however due to the reduction in redundancy the network should be deemed as at risk.

We have contacted the provider of this circuit and will update with any information as we have it.
Mar 14, 23:50 UTC